OpenStudy (anonymous):

Mario obtains a 30/7 balloon mortgage to finance $245,500 at 5.25%. How much principal and interest will he have already paid when his balloon payment is due? $113,875.44 $112,519.78 $127,304.57 $216,988.63

OpenStudy (amistre64):

wouldnt that just be the sum of his payments?

OpenStudy (amistre64):

7*12 payments of .... how do you calculate payments?

OpenStudy (anonymous):

84

OpenStudy (amistre64):

84 payments, yes ..... how do you calculate a payment amount?

OpenStudy (anonymous):

Idk on that part .

OpenStudy (amistre64):

ive responded to enough of your posts that i know that you have some idea, or formula, or attempt

OpenStudy (dumbcow):

use a finance calculator like a TI BA

OpenStudy (amistre64):

ive got enough practice determining a payment amount, i need to know what your process is :)

OpenStudy (dumbcow):

@amistre64 , ive noticed you use a different formula than what i use

OpenStudy (amistre64):

i do :) it my own creation

OpenStudy (dumbcow):

:)

OpenStudy (amistre64):

it simplifies to the textbook formula, but i find my version more intuitive is all

OpenStudy (anonymous):

I just don't understand your formula that's all -.-

OpenStudy (amistre64):

then lets use your formula; they are elementary the same

OpenStudy (anonymous):

B=B[((1+i)^(n*t)-(1+i)^(m))/((1+i)^(n*t)-(1))

OpenStudy (anonymous):

I just dont know what m = ?

OpenStudy (amistre64):

m looks to be the number of months

OpenStudy (amistre64):

for me: remianing balance = loan compounded - payments(geo sum) 0 = loan compounded - payments(geo sum) loan compounded = payments(geo sum) loan compounded/(geo sum) = payments 245500k^(12*30)(1-k)/(1-k^(12*30)), k=1+.0525/12 abd we have 84 payments 84*245500k^(12*30)(1-k)/(1-k^(12*30)), k=1+.0525/12 http://www.wolframalpha.com/input/?i=84*245500k%5E%2812*30%29%281-k%29%2F%281-k%5E%2812*30%29%29%2C+k%3D1%2B.0525%2F12

OpenStudy (anonymous):

P=245500 r=0.0525 n=12 i=r/n=0.0525/12=0.004375 t=30 m=7 Am I right ? :D

OpenStudy (amistre64):

\[B\frac{(1+i)^{(n*t)}-(1+i)^m}{(1+i)^{(n*t)}-1}\] \[254500\frac{(1+.0525/12)^{(12*30)}-(1+.0525/12)^{(84)}}{(1+.0525/12)^{(12*30)}-1}\]maybe

OpenStudy (amistre64):

well, 245500

OpenStudy (anonymous):

ohhh so m=84 :D

OpenStudy (amistre64):

maybe ..... when i use 84 with that i end up with the 216 opt

OpenStudy (amistre64):

i get the 113 with my thought; sum of the payments is how much youve paid in principal and interest

OpenStudy (anonymous):

so 216 is the correct answer ?

OpenStudy (amistre64):

i believe the correct answer is the sum of all the payments youve made; which is what i get as 113, i cant make sense of the formula you presented unless it might be fore the balloon payment

OpenStudy (anonymous):

ok thank you !

OpenStudy (amistre64):

216 989 is the balloon payment, or whats owed on it 113 875 is how much youve already spent

OpenStudy (anonymous):

So, does that mean the answer is 216,989? I'm working the same problem and I'm a little confused. When I used the formula given to me, I got 216,988.63 (which can be rounded to 216,989). I'm just not sure if its the right answer.

OpenStudy (amistre64):

the question is asking: how much has already been paid up when the ballon payment is due. 216k is what is owed, it is the balloon payment. 114k is what has already been spent in principal and interest.